Grade 11Mechanics

The acceleration a of body moving with initial velocity u changes with distance x as a=k^2 root x where k is a positive constant. The distance travelled by the body when its velocity becomes 2u is

Anonymous
9 Years agoGrade 11

2 Answers

Ritayan Chakrabarti
8 Years ago
The answer is (3u/2k)⅔ this u can get by the integration of the givenformula. U see, a=k²√x so V= 2/3k²yx²/³ where y=dt/dx
Ritayan Chakrabarti
8 Years ago
As fr ritayan, the y is considerdd to be equal to 1 fr constructive purposee.